OpenStudy (anonymous):

Solve the equation Quantity x minus 6 all over 2 = 3 for x.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hellp

OpenStudy (anonymous):

!!!! :,(

OpenStudy (anonymous):

lol "all over 2 "

OpenStudy (anonymous):

need the ans. again?

OpenStudy (anonymous):

i got it here

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol... im pretty sure ur using apex... right?

OpenStudy (anonymous):

well... that ans. is 6

OpenStudy (anonymous):

yes! thanks i have one more question to answer

OpenStudy (anonymous):

sure... what?

OpenStudy (anonymous):

Solve: x − 3 = −12

OpenStudy (anonymous):

lol what is 9-3?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

it's 6... but backwards its...?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

-9-3?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

lol noo... its the answer ur looking ofr

OpenStudy (anonymous):

huh ?

OpenStudy (anonymous):

-9-3=-12

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

>.< thanks :*

OpenStudy (anonymous):

what grade are u in?!

OpenStudy (anonymous):

1st

OpenStudy (anonymous):

wait thats not one of the answers and im a junior i failed algebra .....

OpenStudy (anonymous):

lol and im a soon-to-be senior

OpenStudy (anonymous):

lmaoooooo a junior in high scholl still in algebra 1 dang drop out :/

OpenStudy (anonymous):

cant wait fo da treats :D

OpenStudy (anonymous):

lol ur going to have a hard time findin a job :P

OpenStudy (anonymous):

im not a drop out ? o.o i failed algebra when i was in ninth im still in school i was always bad at math

OpenStudy (anonymous):

im talking to u @NikoTheBaller

OpenStudy (anonymous):

Im not a drop out lol im in 11th Algebra 2

OpenStudy (anonymous):

and Solve: x − 3 = −12 its not - 12 o.o

OpenStudy (anonymous):

@Chrisgie

OpenStudy (anonymous):

Your not gonna pass geometry no chance if you have trouble with this sorry to say......... @Notamathgenius01

OpenStudy (anonymous):

hey! @Notamathgenius01 ... the answer is -9

OpenStudy (anonymous):

thanks and @NikoTheBaller i passed geometry

OpenStudy (anonymous):

How did you pass geometry but fail algebra 1 ?.????

OpenStudy (anonymous):

must've retook it during the school yr

OpenStudy (anonymous):

kinda when i flunked 1 sem of pre-alg

OpenStudy (anonymous):

but im going fine now :)

OpenStudy (anonymous):

geometry was easy for me and dudes or watever you guys are g2g im a chick btw hahha js

OpenStudy (anonymous):

lol didn't kno that

OpenStudy (anonymous):

Okay and..............

OpenStudy (anonymous):

and im a guy lol